OpenStudy (anonymous):

A 228 g metal bar requires 5.20 kJ to change its temperature from 11.8°C to 90.2°C. What is the specific heat of the metal?

OpenStudy (john_es):

Same formula, \[Q=m\cdot c\cdot\Delta T\] But now, you need c, \[c=\frac{Q}{m\cdot \cdot\Delta T}\]

OpenStudy (john_es):

I recommend you to use SI units, so put Q in Joules.

OpenStudy (anonymous):

ok and how do you do significant figures

OpenStudy (john_es):

significant figures?

OpenStudy (anonymous):

yes its something my teacher told us. it determines where you put the decimal

OpenStudy (john_es):

Ah, ok. You must use the following rule. Always put the minimum number of significan figures when operate with numbers. For example, the number 9.28 has 3 significant figures. Then number 0.3 has one significant figure (left zeros not count). If I operate with them, \[0.03\cdot9.28=0.2784\] But I have only one signficant figure, so I must put, \[0.3\] as the final result.

OpenStudy (john_es):

A more detailed explanation can be found here for more general cases, http://en.wikipedia.org/wiki/Significant_figures

OpenStudy (anonymous):

so what do you base the number of sigfigs in the equation on

OpenStudy (anonymous):

like what would be the answer using sigfigs for A 228 g metal bar requires 5.20 kJ to change its temperature from 11.8°C to 90.2°C. What is the specific heat of the metal?

OpenStudy (john_es):

All numbers you use has three digits, so it must be three significant figures.

OpenStudy (john_es):

\[5.20/(228\cdot78.4)=2.91\cdot10^{-4}\ kJ/(g\cdot K)=2.91\cdot10^{-1}\ J/(g\cdot K)\]